OpenStudy (anonymous):

Find the next term in this arithmetic sequence: -3, -5, -7

OpenStudy (anonymous):

-9 ?

sam (.sam.):

If you want to express it in calculations, \[a_1=-3~~,~~d=-5-(-3)=-2\] \[T_4=a_{1}+(n-1)d\] \[T_4=-3+(3)(-2)=-9\]